OpenStudy (sphott51):

(-1 +6i) + (-4 + 2i)

OpenStudy (anonymous):

the answer is (1+2i)

OpenStudy (sphott51):

Not one of the answer choices

OpenStudy (anonymous):

what are the answer choices

OpenStudy (sphott51):

a: 5-8i b:5-2i c:-5+8i d:3i

OpenStudy (tkhunny):

It would help if you wrote the original problem correctly. It would also help if you showed YOUR work.

OpenStudy (anonymous):

combine like terms that is all

OpenStudy (sphott51):

4-2i -24i +12i^2 4 -26i -12 = -8 -26i ...?

OpenStudy (usukidoll):

this is addition right? or is it multiplication? oh man I agree with tkhunny.

OpenStudy (anonymous):

lol at the risk of repeating myself, combine like terms \[-1+6x-4+2x\]

OpenStudy (usukidoll):

I think...just a suggestion that if the original poster is typing the equations or questions wrong, we shouldn't go further unless there is more information.

OpenStudy (sphott51):

I fixed it.. happy now?

OpenStudy (usukidoll):

much better.. now do what satellite said.

OpenStudy (sphott51):

This is an imaginary number...

OpenStudy (anonymous):

then imagine that \(x\) is \(i\) it is the same process

OpenStudy (sphott51):

isn't i^2 = -1 ?

OpenStudy (anonymous):

yes, but there is no \(i^2\) in your question think you might be confusing addition and multiplication

OpenStudy (usukidoll):

true, but you're just adding and combining like terms for this one

OpenStudy (usukidoll):

a typical problem for multiplication would look like this (a+bi)(a-bi) but you don't have that

OpenStudy (sphott51):

oh........ thanks

OpenStudy (usukidoll):

for this addition problem your final answer should look like one of these four a+bi,a-bi,ai+b,ai-b

OpenStudy (usukidoll):

the order doesn't matter that much

OpenStudy (sphott51):

yeah i got -5 +8i , it's one of the answer choices.

OpenStudy (usukidoll):

that's correct